OpenStudy (anonymous):

what number is 73% of 215 A.215=0.73n B.n=0.73 x 215 C.0.73=n x 215 D.215=73n what percent of 94 is 23? A.n/94=23/100 B.23/94=n/100 C.94/n=23/100 D.23/94=100/n

OpenStudy (anonymous):

73% = \(\dfrac{73}{100}\)

OpenStudy (anonymous):

All you have to do is multiply that by the of number...

OpenStudy (anonymous):

A percent would mean you's always divide a number with this sign % by 100.

OpenStudy (jidnyesha):

\[\frac{ 73 }{ 100}*215=n\] n=.73*215

OpenStudy (anonymous):

i am not good at math can u help me ??

OpenStudy (jidnyesha):

23/94*100=n

OpenStudy (jidnyesha):

73 percent means 73/100 "of" is taken as multplication so 73/100*215=n

OpenStudy (anonymous):

It's best to not give direct answers @Jidnyesha.

OpenStudy (jidnyesha):

oh i thought i gave the steps too

OpenStudy (anonymous):

i need the answer for both asap plz

OpenStudy (jidnyesha):

anyways srry ill give the explaination too actually im new here

OpenStudy (anonymous):

No worries- welcome to Open Study :)

OpenStudy (anonymous):

" is " means equals, and " of " means multiply

OpenStudy (jidnyesha):

for the second one when it is said to express a certain number as a % of the other we multiply the fraction obtained by 100 so when it is said to express 23 as % of 95 the fraction will be 23/95 thus 23/95*100 will be the answer